# A systematic search for AGN obscuration variability in the Chandra archive

### v1

@@ 5.1 Number of observations
This is suggesting the same trend found here, and perhaps the reason most of their sources are not variable is simply due to the limited amount of observations for each source. We believe this is a fair comparison despite the fact that Laha et al. (2020) is directly measuring and uses stricter criteria for defining a ‘variable’ source because the trend is still that variable determinations tend to come from groups with more observations, while ‘non-variable’ sources tend to have fewer observations, agreeing with intuition.
@@
It should also be noted that the trend in this analysis may simply be due to false positives. As discussed in Cox et al. (2023), it is possible to observe measurable differences in hardness ratios between spectra simulated with the exact same conditions. As the number of observations increases, the likelihood of measuring such a false positive increases. However, assuming FPR=0.4, for any given source, only 2.5 observation pairs would need to be flagged for a 90 % probability that at least one pair (and therefore the source) is truly variable. Of the 43 variable sources, 21 have at least 3 variable pairs (8 sources only have two observations and cannot have 3 variable pairs). However, this explanation cannot be used to motivate the trend in Laha et al. (2020) since their results are based on a detailed spectral analysis. In any case, the observed fraction should still be regarded as a lower limit that will likely increase as more data becomes available.
5.2 HR1 vs HR2 and other sources of variability
@@
Across the 43 flagged sources, there were 378 observation pairs that were flagged. Of those, 240 were flagged in only HR1, 85 were flagged in only HR2, and 53 were flagged in both HR1 and HR2. The values in both HR1 and HR2 are shown in the left panel of Figure 4. As discussed in Cox et al. (2023), HR1 is most sensitive to variability in the range cm-2 while HR2 is most sensitive to changes in the range cm-2. These results indicate that much of the variability could be coming from only moderately obscured AGN.
@@
Of course, a hardness ratio will detect any change in the spectral shape including those not due to changes in (e.g. photon index, reflection, non-AGN related changes). However, if the photon index changes, one would expect to detect the change in both HR1 and HR2. Only 53 out of 378 flagged events were detected in both ratios, so many flagged events cannot be attributed to a change in photon index. We also investigated how large a change in photon index would be required before our method would detect it above the expected false positive rate. We found that our method is not affected at all for observation pairs with average counts when photon index is varied between and ( % of the total sample). For higher average counts (), large changes in photon index can potentially increase the false positive rate of our detection. At cm-2, the change in photon index must be . At higher obscuration levels, changes in photon index would have to be even larger (and physically unlikely). More specifically, at cm-2, and at cm-2, . Only % of the observation pairs have average counts so it is unlikely changes in photon index are contributing significantly to our statistics.
@@
Furthermore, assuming the reflection component comes primarily from the parsec-scale torus, we would not expect it to change significantly in less than 104 days (Marchesi et al., 2022; Torres-Albà et al., 2023, 2025). Finally, non-AGN emission is typically most prominent at energies keV which is excluded from both HR1 and HR2. Therefore, we believe it is reasonable to expect most of these flagged observation pairs are due to changes in . We will perform detailed spectral modeling in a future work to confirm this and determine the magnitude of the suspected variability.
5.3 Data quality bias
@@
Given the nature of the flagging mechanism and the dependence on the size of the HR errors, this method may be prone to simply select high count observations while ignoring all the low count observations, regardless of true variability. To test this, we compare the count levels of the flagged observation pairs and the unflagged observation pairs. The right panel of Figure 4 shows the distributions of the average counts between the two observations in each flagged pair and each unflagged pair. As expected, the flagged observation pairs tend to have slightly higher counts since at a given level of variability, higher quality data will be more likely to detect it. For the unflagged observation pairs, 95 % of them have between 66 and 609 counts. For the flagged observation pairs, 95 % of them have between 94 and 2366 counts. There are many high count observations that remain unflagged and many low count observations that are flagged. Therefore, we are confident that this method is not simply selecting high count observations while ignoring low count observations. If the observation pair is flagged, there is likely a significant difference in the spectral shape regardless of data quality. On the other hand, if the observation pair is not flagged, there is likely no observable difference in spectral shape, although this is less certain for low quality data. We are not overly concerned about missing potential false negatives since the primary goal of this work is to provide a manageable sample of sources that are most likely variable in and to provide a lower limit on the -variable fraction in the local universe for the first time. This lower limit remains valid no matter how high the false negative rate is.
5.4 Time intervals between observations
@@
There are 1,455 observation pairs in the sample. The statistics of each pair as a function of the time separation between them are plotted in the left panel of Figure 5. The flagging threshold of is shown as a black dotted line. Previous studies have indicated a potential dependence between variability and timescale (e.g., Torres-Albà et al., 2023; Pizzetti et al., 2025; Torres-Albà et al., 2025). That is, longer timescales tend to have a higher likelihood of showing variability. The idea is that variability on short timescales is likely due to broad line region clouds passing in and out of the line of sight, whereas variability on longer timescales is likely due to material further away from the corona, such as the dusty torus. However, lacking a continuous monitoring campaign to measure ingresses and egresses of eclipses on all timescales, we are forced to make conclusions on population statistics.
@@
To test this idea, we split our data into short timescales (100 days) and long timescales (100 days) to see if there is a difference in the fraction of observations flagged as variable. The distributions for the short and long timescale populations are shown in the right panel of Figure 5. The flagging threshold of is shown as the vertical dashed line. At long timescales, % of the observation pairs were flagged as variable while at short timescales, % were flagged. This is in agreement with the results from Torres-Albà et al. (2023) and Pizzetti et al. (2025). A Kolmogorov-Smirnov (KS) test indicates a significant difference between the distributions with p-value for a two-sided test as well as the alternative hypothesis that the CDF of the short timescales being greater than the long timescale (i.e. long timescales have higher values).
@@
5.5 Seyfert 1 vs. Seyfert 2 galaxies
@@
Finally, in Seyfert 1 galaxies, the obscuration would likely be caused by broad line region clouds (e.g., Elvis et al., 2004; Puccetti et al., 2007; Maiolino et al., 2010). Since these are closer to the SMBH and have faster orbits, there is variability present at all timescales compared to Seyfert 2 galaxies where obscuration is likely dominated by the torus and therefore detecting variability would be restricted to longer timescales (e.g., Torres-Albà et al., 2025). Figure 5 shows a hint of this. The median value for Seyfert 1 galaxies is higher and there is a higher percentage of observation pairs flagged as variable at all timescales. However, this difference is very slight and could still be explained by the fact that the Seyfert 1 observation pairs have higher quality data.
5.6 PGC 94626
PGC 94626 has been classified as both a type 1 (Lin et al., 2012) and a type 2 (Khamitov et al., 2023) galaxy, however, we see no evidence of type 2 behavior in X-ray or optical emission. The X-ray spectrum is not heavily absorbed and there are broad lines present in the optical spectrum. We checked the [OIII]/H line ratio and found it to be consistent with a Seyfert 1.2 (Osterbrock, 1977; Whittle, 1992). For these reasons, we opt for the Seyfert 1 classification, consistent with the classification in the Milliquas Catalog. We also checked available NuSTAR and XMM-Newton data, observation IDs 70660003002 (July 15, 2020) and 0097820101 (June 26, 2000), respectively. PGC 94626 is visible in the NuSTAR data, however, there is a significant amount of background with it being so close to Abell 1795. We obtain useful data between and 10 keV and the spectrum shows very low obscuration with a photon index of . The XMM-Newton image is less significantly background dominated, however there is a CCD line passing by the source at a distance of ” in the PN image and ” in the MOS1 and MOS2 images. We therefore extract spectra using circular regions with radii of 20” and 10” for PN and MOS, respectively. The spectrum is again clearly unabsorbed with a photon index
@@ 5.6 PGC 94626
The HR curve shown in Figure 6 only shows small fluctuations, with most measurements of HR1 falling between and , and most of the HR2 measurements between and . Since the source is mostly unabsorbed, the data with energies keV does not change much, so the sensitivity of our HR bands are limited. However, many observation pairs are flagged (433 in HR1, 237 in HR2, and 51 in both) and the pair studied here differed in by a factor of . Therefore, this method is still capable of detecting large changes in even when the source has cm-2. However, we note that using the energies keV would likely produce a higher true positive rate for such a low-obscuration source, if not for the sensitivity degradation in Chandra.
@@ 6 Summary
In this work, we have applied a hardness ratio variability analysis to 79 sources (41 Seyfert 1 and 38 Seyfert 2 galaxies) in the local universe () with at least two Chandra observations (§2). This analysis was specifically tailored for detecting changes in in moderately obscured AGN and has an expected true positive rate of % and an expected false positive rate of %. We find that % of the total sample displays variability in the hardness ratios when using as the flagging threshold. Using the flagging threshold , which should have almost no false positive detections, the observed fraction is %, however, this number goes up to % when considering sources with at least 5 observations (§4).
@@
We find that sources with more observations are more likely to be flagged as variable (§5.1). All 11 sources with at least 10 observations were flagged as variable while 8/31 sources with only 2 observations were flagged. Results from previous studies (e.g. Laha et al., 2020) have suggested the same trend. This could indicate that all sources are variable, and whether or not they are detected as such depends primarily on how often we observe them. Therefore, the variability fraction presented here should be taken as a lower limit, and we expect it to increase as more data becomes available on previously observed sources. Considering the highest threshold and sources with a reasonable number of observations (5–10), the fraction increases to % (see inset in Figure 2).
@@ 7 Acknowledgments
I.C. acknowledges support under grant AR4-25009X. The scientific results reported in this article are based on observations made by the X-ray observatories Chandra. We have made use of the NASA/IPAC Extragalactic Database (NED), which is operated by the Jet Propulsion Laboratory, California Institute of Technology under contract with NASA. Furthermore, this research has made use of the VizieR catalogue access tool, CDS, Strasbourg, France https://doi.org/10.26093/cds/vizier. The original description of the VizieR service was published in Ochsenbein et al. (2000). We have made extensive use of the packages dplyr (Wickham et al., 2023) and ggplot2 (Wickham, 2016) within the R programming language (R Core Team, 2021). We also acknowledge use of the NumPy (Harris et al., 2020) and Pandas (McKinney et al., 2010) packages in the Python3 (Van Rossum & Drake, 2009) programming language.

### v2

@@ 5.1 Impact of the number of observations on the observed variable fraction
This is suggesting the same trend found here, and perhaps the reason most of their sources are not variable is simply due to the limited amount of observations for each source. We believe this is a fair comparison despite the fact that Laha2020 is directly measuring and uses stricter criteria for defining a ‘variable’ source because the trend is still that variable determinations tend to come from groups with more observations, while ‘non-variable’ sources tend to have fewer observations, agreeing with intuition.
However, this trend could potentially be spurious if sources known to be variable are being targeted more often due to their variability. Another option is that a large number of observations per source guarantees a false positive. We consider both of these below.
Targeting variable sources. It could be that known variable sources have more observations because they are targeted more frequently to analyze their variable nature. However, this does not appear to be the case here based on a search of the literature and proposal abstracts. Only two sources, ESO 253-3 (with 11 observations used) and PGC 214276 (with 12 observations used), were monitored specifically for expected variability. In the case of ESO 253-3, there is a suspected periodic tidal disruption event (e.g. ID: 21708757, P.I. Anna Payne; ID: 23700492, P.I. Katie Auchettl; ID: 24700009, P.I. Norbert Schartel; ID: 26700325, P.I. Katie Auchettl), and in the case of PGC 214276, they were hoping to observe a SMBH merger (ID: 23708834, P.I. Ning Jiang). In both cases, the accretion properties would be changing rapidly, potentially resulting in a large change in photon index which may be the cause of our variability detection. Removing these two sources from the list of variable sources would result in a variable fraction of %. To our knowledge, no other sources identified as variable in this study were targeted for follow-up observations based on known variability. Therefore the trend observed here is not significantly affected by this bias.
@@
False positives. It should be noted that the trend in this analysis may simply be due to false positives. As discussed in Cox2023, it is possible to observe measurable differences in hardness ratios between spectra simulated with the exact same conditions. As the number of observations increases, the likelihood of measuring such a false positive increases. However, assuming FPR=0.4, for any given source, only 2.5 observation pairs would need to be flagged for a 90 % probability that at least one pair (and therefore the source) is truly variable. Of the 43 variable sources, 21 have at least 3 variable pairs (8 sources only have two observations and cannot have 3 variable pairs). However, this explanation cannot be used to motivate the trend in Laha2020 since their results are based on a detailed spectral analysis.
We therefore believe that this trend should be considered as a true bias when determining whether sources are non-variable. Using the inset of Figure 2 as a guide, the variable fraction appears to level off between 5 and 10 observations555Since there are only 9 sources with more than 10 observations, the measured fraction begins to lose meaning due to the small number of sources.. As such, judgement should be withheld on non-variable sources with fewer than 5–10 observations as it is likely they have not been observed enough to discover their variability. On the other hand, a source that remains non-variable after 5–10 observations is potentially interesting. For example, this is the case for NGC 6300 which was shown by Sengupta25 to lack significant variability across 10 observations taken over 13 years.
Given these considerations, we believe the best estimate of the variable fraction comes from applying the most conservative threshold to eliminate the possibility of false positives, and only considering sources with 5–10 observations. The fraction obtained in this way is 50 %. However, we caution that this should still be considered a lower limit.
5.2 Other sources of spectral variability
@@
Across the 43 flagged sources, there were 378 observation pairs that were flagged. Of those, 240 were flagged in only HR1, 85 were flagged in only HR2, and 53 were flagged in both HR1 and HR2. The values in both HR1 and HR2 are shown in the left panel of Figure 5. As discussed in Cox2023, HR1 is most sensitive to variability in the range cm-2 while HR2 is most sensitive to changes in the range cm-2. These results indicate that much of the variability could be coming from only moderately obscured AGN.
Of course, a hardness ratio will detect any change in the spectral shape including those not due to changes in (e.g. photon index, reflection, non-AGN related changes).
@@
Photon Index. If the photon index changes, one would expect to detect the change in both HR1 and HR2. Only 53 out of 378 flagged events were detected in both ratios, so many flagged events cannot be attributed to a change in photon index. We also investigated how large a change in photon index would be required before our method would detect it above the expected false positive rate. We found that our method is not affected at all for observation pairs with average counts when photon index is varied between and ( % of the total sample). For higher average counts (), large changes in photon index can potentially increase the false positive rate of our detection. At cm-2, the change in photon index must be . At higher obscuration levels, changes in photon index would have to be even larger (and physically unlikely). More specifically, at cm-2, and at cm-2, . Only % of the observation pairs have average counts so it is unlikely changes in photon index are contributing significantly to our statistics. Figure 4 shows examples of the change in hardness ratios for different photon index values and data quality. These plots also show that we are not sensitive to flux variability, as expected.
It is also worth noting that we have no reason to believe there could be such extreme variability in photon index occurring in this sample. For example, in previous studies of variability for almost 30 type 2 sources, none required a variable photon index to achieve a good fit, although many required flux variability (pizzetti_multi-epoch_2022; NTA23; Pizzetti2025; NTA25; Sengupta25).
Reflection. Assuming the reflection component comes primarily from the parsec-scale torus, we would not expect it to change significantly in less than 104 days (marchesi_compton-thick_2022; NTA23; NTA25).
Non-AGN emission. Non-AGN emission is typically most prominent at energies keV which is excluded from both HR1 and HR2. As with torus reflection, we also would not expect much variability from these regions on the timescales observed.
@@
Therefore, we believe it is reasonable to expect most of these flagged observation pairs are due to changes in . We will perform detailed spectral modeling in a future work to confirm this and determine the magnitude of the suspected variability.
5.3 Impact of data quality on variability detection
@@
Given the nature of the flagging mechanism and the dependence on the size of the HR errors, this method may be prone to simply select high count observations while ignoring all the low count observations, regardless of true variability. To test this, we compare the count levels of the flagged observation pairs and the unflagged observation pairs. The right panel of Figure 5 shows the distributions of the average counts between the two observations in each flagged pair and each unflagged pair. As expected, the flagged observation pairs tend to have slightly higher counts since at a given level of variability, higher quality data will be more likely to detect it. For the unflagged observation pairs, 95 % of them have between 66 and 609 counts. For the flagged observation pairs, 95 % of them have between 94 and 2366 counts. However, our detection method is still sensitive in the low-count regime as variability is detected in many observation pairs with only counts. Furthermore, there are many observation pairs with counts that remain non-detections. Therefore, we are confident that this method is not simply selecting high count observations while ignoring low count observations. If the observation pair is flagged, there is likely a significant difference in the spectral shape regardless of data quality. On the other hand, if the observation pair is not flagged, there is likely no observable difference in spectral shape, although this is less certain for low quality data. We are not overly concerned about false negatives since the primary goal of this work is to provide a lower limit on the -variable fraction in the local universe for the first time.
Importantly, Cox2023 showed that the false positive rate appears to be completely independent of data quality based off of analysis of simulated spectra. This is due to the fact that low quality spectra result in large scatter in the HR measurements of spectra with a given , but also increased errors on the measurements. These two effects offset each other in the final prediction, resulting in the same false positive rate for the entire range between 100 counts and 20,000 counts (see Figure 15 in Cox2023).
Additionally, we tested that our cut requiring counts truly allowed us to apply the classical error propagation assuming Gaussian statistics, as opposed to the more rigorous calculation of park_bayesian_2006. We calculated each of our hardness ratios using the BEHR code presented in park_bayesian_2006, and found that only % of error calculations were significantly different from those obtained with the classical method. More importantly, only 7 observations have classical calculations that underestimate the error relative to the BEHR method. In these cases, we checked the variability prediction given by the BEHR hardness ratios and found that our conclusions do not change. In all other cases, the classical errors overestimate the size of the errors relative to BEHR, making our lower limit even more robust.
5.4 Correlation of variable fraction with the time interval between observations
@@
There are 1,455 observation pairs in the sample. The statistics of each pair as a function of the time separation between them are plotted in the left panel of Figure 6. The flagging threshold of is shown as a black dotted line. Previous studies have indicated a potential dependence between variability and timescale (e.g., NTA23; Pizzetti2025; NTA25). That is, longer timescales tend to have a higher likelihood of showing variability. The idea is that variability on short timescales is likely due to broad line region clouds passing in and out of the line of sight, whereas variability on longer timescales is likely due to material further away from the corona, such as the dusty torus. However, lacking a continuous monitoring campaign to measure ingresses and egresses of eclipses on all timescales, we are forced to make conclusions on population statistics.
@@
To test this idea, we split our data into short timescales (100 days) and long timescales (100 days) to see if there is a difference in the fraction of observations flagged as variable. The distributions for the short and long timescale populations are shown in the right panel of Figure 6. The flagging threshold of is shown as the vertical dashed line. At long timescales, % of the observation pairs were flagged as variable while at short timescales, % were flagged. This is in agreement with the results from NTA23 and Pizzetti2025. A Kolmogorov-Smirnov (KS) test indicates a significant difference between the distributions with p-value for a two-sided test as well as the alternative hypothesis that the CDF of the short timescales being greater than the long timescale (i.e. long timescales have higher values).
@@
5.5 Difference in variable fraction between Seyfert 1 and Seyfert 2 galaxies
@@
Finally, in Seyfert 1 galaxies, the obscuration would likely be caused by broad line region clouds (e.g., Elvis2004; Puccetti2007; Maiolino2010). Since these are closer to the SMBH and have faster orbits, there is variability present at all timescales compared to Seyfert 2 galaxies where obscuration is likely dominated by the torus and therefore detecting variability would be restricted to longer timescales (e.g., NTA25). Figure 6 shows a hint of this. The median value for Seyfert 1 galaxies is higher and there is a higher percentage of observation pairs flagged as variable at all timescales. However, this difference is very slight and could still be explained by the fact that the Seyfert 1 observation pairs have higher quality data.
5.6 The case of PGC 94626
PGC 94626 has been classified as both a type 1 (Lin12) and a type 2 (Khamitov23) galaxy. However, we see no evidence of type 2 behavior in X-ray or optical emission. The X-ray spectrum is not heavily absorbed and there are broad lines present in the optical spectrum. We checked666The SDSS spectrum can be found at https://skyserver.sdss.org/dr19/VisualTools/explore/summary?sid=2380220362467076096. the [OIII]/H line ratio and found it to be consistent with a Seyfert 1.2 (Osterbrock77; Whittle92). For these reasons, we opt for the Seyfert 1 classification, consistent with the classification in the Milliquas Catalog. We also checked available NuSTAR and XMM-Newton data, observation IDs 70660003002 (July 15, 2020) and 0097820101 (June 26, 2000), respectively. PGC 94626 is visible in the NuSTAR data, however, there is a significant amount of background with it being so close to Abell 1795. We obtain useful data between and 10 keV and the spectrum shows very low obscuration with a photon index of . The XMM-Newton image is less significantly background dominated, however there is a CCD line passing by the source at a distance of ” in the PN image and ” in the MOS1 and MOS2 images. We therefore extract spectra using circular regions with radii of 20” and 10” for PN and MOS, respectively. The spectrum is again clearly unabsorbed with a photon index
@@ 5.6 The case of PGC 94626
The HR curve shown in Figure 7 only shows small fluctuations, with most measurements of HR1 falling between and , and most of the HR2 measurements between and . Since the source is mostly unabsorbed, the data with energies keV does not change much, so the sensitivity of our HR bands is limited. However, many observation pairs are flagged (433 in HR1, 237 in HR2, and 51 in both) and the pair studied here differed in by a factor of . Therefore, this method is still capable of detecting large changes in even when the source has cm-2. However, we note that using the energies keV would likely produce a higher true positive rate for such a low-obscuration source, if not for the sensitivity degradation in Chandra.
@@ 6 Summary
In this work, we have applied a hardness ratio variability analysis to 79 sources (41 Seyfert 1 and 38 Seyfert 2 galaxies) in the local universe () with at least two Chandra observations (§2). This analysis was specifically tailored for detecting changes in in moderately obscured AGN and has an expected true positive rate of % and an expected false positive rate of %. We find that % of the total sample displays variability in the hardness ratios when using as the flagging threshold. Using the flagging threshold , which should have almost no false positive detections, the observed fraction is %, however, this number goes up to % when considering sources with at least 5 observations (see §4 and §5.1).
@@
We find that sources with more observations are more likely to be flagged as variable (§5.1). All 11 sources with at least 10 observations were flagged as variable while 8/31 sources with only 2 observations were flagged. Results from previous studies (e.g. Laha2020) have suggested the same trend. This could indicate that all sources are variable, and whether or not they are detected as such depends primarily on how often we observe them. Therefore, the variability fraction presented here should be taken as a lower limit, and we expect it to increase as more data becomes available on previously observed sources. Considering the highest threshold and sources with a reasonable number of observations (5–10), the fraction increases to % (see inset in Figure 2).
@@ 7 Acknowledgments
We thank the anonymous referee for their helpful comments which have improved the quality of this manuscript. I.C. acknowledges support under grant AR4-25009X. The scientific results reported in this article are based on observations made by the X-ray observatories Chandra. We have made use of the NASA/IPAC Extragalactic Database (NED), which is operated by the Jet Propulsion Laboratory, California Institute of Technology under contract with NASA. Furthermore, this research has made use of the VizieR catalogue access tool, CDS, Strasbourg, France https://doi.org/10.26093/cds/vizier. The original description of the VizieR service was published in vizier2000. We have made extensive use of the packages dplyr (dplyr) and ggplot2 (ggplot2) within the R programming language (Rcode). We also acknowledge use of the NumPy (numpy) and Pandas (pandas) packages in the Python3 (python) programming language. This paper employs a list of Chandra datasets, obtained by the Chandra X-ray Observatory, contained in the Chandra Data Collection (CDC) at https://doi.org/10.25574/cdc.461 (catalog doi:10.25574).